Two loudspeakers placed 6.0 m apart are driven in phase by an audio oscillator, whose frequency range is 711 Hz to 1124 Hz. A point P is located 5.1 m from one loudspeaker and 3.6 m from the other. The speed of sound is 344 m/s. The frequency produced by the oscillator, for which destructive interference occurs at point P, in SI units, is closest to: Two loudspeakers placed 6.0 m apart are driven in phase by an audio oscillator, whose frequency range is 711 Hz to 1124 Hz. A point P is located 5.1 m from one loudspeaker and 3.6 m from the other. The speed of sound is 344 m/s. The frequency produced by the oscillator, for which destructive interference occurs at point P, in SI units, is closest to:

a. 1055

b. 1032

c. 1078

d. 1009

e. 1101
Answer #1

  Constructive interference: |L1-L2|=nλ
nλ= 5.1-3.6= 1.5m

1055 Hz: 1055/(344/1.5)= 4.6λ
1032 Hz: 1032/(344/1.5)= 4.5λ (Integral multiple correct answer)
1078 Hz: 1078/(344/1.5)= 4.7λ
1009 Hz: 1009/(344/1.5)= 4.39λ

1101Hz: 1101/(344/1.5)= 4.8 λ

Ans- b) 1032